# Low-Level and NUMA-Aware Optimization for High-Performance Quantum Simulation

### v1

@@ Abstract
Scalable classical simulation of quantum circuits is crucial for advancing both quantum algorithm development and hardware validation. In this work, we focus on performance enhancements through meticulous low-level tuning on a single-node system, thereby not only advancing the performance of classical quantum simulations but also laying the groundwork for scalable, heterogeneous implementations that may eventually bridge the gap toward noiseless quantum computing. Although similar efforts in low-level tuning have been reported in the literature, such implementations have not been released as open-source software, thereby impeding independent evaluation and further development. We introduce an open‑source, high‑performance extension to the QuEST simulator that brings state‑of‑the‑art low‑level and NUMA optimizations to modern computers. Our approach emphasizes locality-aware computation and incorporates hardware-specific optimizations such as NUMA-aware memory allocation, thread pinning, AVX-512 vectorization, aggressive loop unrolling, and explicit memory prefetching. Experiments demonstrate significant speedups – 5.5–6.5× for single-qubit gate operations, 4.5× for two-qubit gates, 4× for Random Quantum Circuits (RQC), and 1.8× for Quantum Fourier Transform (QFT), demonstrating that rigorous performance tuning can substantially extend the practical simulation capacity of classical quantum simulators on current hardware.
Index Terms:
CPU; quantum simulation; high performance computing; NUMA; SIMD; QuEST; QFT; RQCI Introduction
@@
Classical simulation of quantum circuits is still indispensable for validating algorithms, benchmarking hardware, and debugging control software, while error-corrected quantum computers remain beyond current reach. In the full state‑vector simulation an -qubit state is represented by complex amplitudes, so both the memory footprint (16 bytes per amplitude in double-precision) and the number of floating‑point operations grow exponentially with . Even a modest 35‑qubit circuit already requires 0.5 TiB of memory and billions of complex arithmetic operations, quickly exhausting today’s servers. Scaling beyond a few dozen qubits remains challenging because hardware and software bottlenecks now dominate performance. Memory bandwidth and capacity are primary constraints: even on multi‑socket non‑uniform memory access (NUMA) servers, accessing a ‑element state-vector saturates the memory hierarchy [1]. In classically distributed simulation of quantum computing, the network communication required for qubit gates – whether single- or two-qubit – spanning different nodes represents a huge overhead and often dominates the run time [1, 2]. At the software level, most simulators still under-use low‑level hardware features, leaving appreciable performance on the table and underscoring the need for architecture‑aware refinements.
@@
State-of-the-art/practice. To push the simulation limits, a suite of high‑performance state‑vector simulators has emerged that exploit modern hardware architectures and parallelism [3]. QuEST [4], ProjectQ [5], Intel‑QS (qHiPSTER) [6], Qrack [7], Qulacs [8], Quantum++ [9], qsim [10] and, most recently, PennyLane Lightning [11] are representative state-of-the-art examples – compared in Table I. Among many simulators exploiting multi‑threading (OpenMP), distributed Message Passing Interface (MPI) execution, and GPU off‑loading to enlarge the range of tractable problem sizes, some state-vector simulators accelerate the core operation (applying gate matrices to the state vector) using explicit single-instruction-multiple-data (SIMD) parallelism. By operating on multiple amplitudes in one CPU instruction, SIMD vectorization can significantly speed up state updates [4]. For example, the ProjectQ simulator explicitly supports Intel AVX vector instructions [5]. More recently, Google’s qsim is built with AVX2 and fused multiply-add (FMA) intrinsics, using fused gate kernels to exploit data-level parallelism [10]. Xanadu’s PennyLane Lightning simulator similarly employs explicit intrinsic AVX-512, effectively unrolling loops to operate on 8 or 16 values at a time [11]. These approaches reduce instruction overhead by applying gate operations to multiple state-vector entries concurrently. Academic simulators have reported substantial speedups from SIMD optimizations. The HpQC simulator used AVX2/FMA instructions to maximize SIMD utilization, combined with manual bit-level optimizations [12]. This yielded about an average speedup of 2.20× (GNU compiler) over QuEST in QFT benchmarks on a multi-core CPU. The authors attribute this gain largely to vectorized gate kernels and reduced instruction count. Another effort, PAS [13], introduced a hybrid vectorization scheme alongside other tricks, achieving up to several-fold speedups versus QuEST.
@@
| Simulator | Vectorisation | Manual Unroll | Cache Blocking | OMP | Thread Pinning | NUMA | MPI | GPU |
| ProjectQ | AVX2 (comp.) | ✗ | ✓ | ✓ | — | ✗ | ✗ | ✗ |
| qsim | AVX2 (intrin.) | ✗ | ✓ | ✓ | — | ✗ | ✗ | ✗ |
| Lightning | AVX2/512 (intrin.) | ✓ | ✗ | ✓ | likely | ✗ | ✓ | ✓ |
| HpQC/PAS | AVX2 (intrin.) | ✗ | ✗ | ✓ | — | ✗ | ✓ | ✗ |
| Qulacs | AVX2 (intrin.) | ✗ | ✗ | ✓ | — | ✗ | ✗ | ✓ |
| Qiskit Aer | AVX2 (comp.) | ✗ | ✗ | ✓ | — | ✗ | ✗ | ✓ |
| Intel‑QS | AVX512 (comp.) | ✗ | ✗ | ✓ | likely | ✗ | ✓ | ✗ |
| QuEST | AVX2 (comp.) | ✗ | ✗ | ✓ | — | ✗ | ✓ | ✓ |
| NWQ‑Sim | AVX2 (comp.) | ✗ | ✗ | ✓ | — | ✗ | ✓ | ✓ |
“Aligned Memory”, not included, is largely implicit (most tools use 64‑byte alignment for AVX‑512), while “Thread Pinning” and “NUMA‑Aware allocation” are seldom discussed—Lightning and Intel‑QS are inferred to pin threads via OMP_PROC_BIND.
comp. = compiler auto‑vectorisation; intrin. = handwritten intrinsics.
likely = affinity hinted at in code or docs but not explicitly evaluated in the paper;
✓/✗ = feature explicitly present or absent in publication, documentation or source.
@@
To fully leverage SIMD, simulators ensure proper memory alignment of the state-vector (often 32- or 64-byte aligned for AVX2/512). Although often not explicitly discussed in publications, alignment is a standard practice to avoid penalties when loading vectors into registers. Many simulators also focus on cache-friendly memory access patterns. One common technique is cache blocking – structuring computations to maximize reuse of data in fast caches before moving to slower memory. In quantum circuit terms, this translates to fusing consecutive gate operations so that the state-vector is traversed fewer times. At a higher level, circuits can also be rearranged to minimise communication, effectively cache-blocking at the node level [14, 15]. ProjectQ and Google’s qsim use such gate fusion to improve locality. By multiplying small gate matrices into a larger effective matrix, the simulator can apply a batch of operations with one pass over the state, thus improving cache utilization and reducing memory bandwidth pressure. Beyond fusion, recent research introduced novel data layouts and filters to minimize unnecessary memory traffic. PAS describes a “memory access filtering” method, which avoids touching state-vector regions that are unaffected by a given gate. Likewise, PAS’s use of fast bitwise operations optimizes the index calculations for mapping qubit operations to memory locations, replacing expensive arithmetic or branch-heavy code with efficient bit-manipulation. The HpQC tool also emphasized a custom data structure to exploit spatial locality in the cache. In practice, these low-level memory optimizations yielded measurable speedups: e.g., HpQC reported 1.74x (GNU compiler) speed improvement on RQCs vs. baseline QuEST by cutting down redundant memory accesses.
@@
Today’s Hardware. Modern CPUs offer dozens of cores, and virtually all state-vector simulators use multithreading to split the amplitude space across cores. OpenMP is commonly employed to parallelize the innermost loops of state updates. QuEST and others leverage multi-core parallelism by dividing the state-vector into chunks processed concurrently by threads. This yields near-linear speedup with core count [4] until memory bandwidth becomes the bottleneck [16]. However, multithreading on large shared-memory systems introduces challenges related to thread affinity and load balancing. If the threads are not carefully pinned or if the work is uneven, performance can suffer. Notably, thread pinning (assigning each thread to a specific core) is often needed to prevent Operating System (OS) scheduler overhead or migration between cores. Additionally, on multi-socket NUMA nodes, pinning helps keep threads on the socket where their data resides. While most quantum simulator papers do not explicitly detail their thread affinity policies, the need is recognized in the community [17].
@@
On NUMA architectures, memory is physically partitioned, and accessing remote memory across the interconnect (e.g., QPI/UPI) happens transparently, i.e., without explicit communication, but incurs higher latency and lower bandwidth. It is therefore crucial for large state-vector simulations to be NUMA-aware. A straightforward strategy is to partition the state-vector across NUMA nodes and pin threads such that each NUMA domain primarily works on its portion of the state. In an ideal scenario, each thread only accesses the segment of the state-vector that resides in its local node’s memory [17]. This avoids the costly scenario where a thread continuously fetches data located in another socket’s RAM. When thread affinity and memory placement are aligned, memory bandwidth scales with the number of NUMA nodes, instead of being bottlenecked by slower cross-socket memory access [18]. Despite its importance, NUMA-aware optimization has not always been a focal point in quantum simulator research. Many open-source simulators rely on the OS’s default first-touch policy to allocate memory locally, but do not report additional NUMA-specific techniques in their publications. We found little explicit analysis of NUMA effects in simulator papers, indicating a gap that our work aims to fill. A recent IBM patent [17] describes how a naive parallelization of a CNOT gate can lead to one socket’s thread constantly accessing another socket’s memory (an imbalanced situation) and proposes reordering qubit indices or replicating data to balance the memory accesses. The fact that such patent-level solutions exist suggests that while NUMA issues are known, they have not been thoroughly addressed in open literature.
Contributions. Motivated by these observations, and building on the strong foundation of QuEST v3.7111In preparing this draft, we noted that QuEST v4 was released, introducing several software architectural enhancements – including revised data structures that partially overlap with our improvements – but it does not incorporate explicit SIMD vectorization or NUMA-aware optimizations as presented in this work. We provide further details in Section V., we introduce a suite of single‑node enhancements intended as a springboard for efficient multi‑node and GPU scaling. Our extension offers tunable modules: (i) a redesigned data layout paired with hand‑written AVX‑512 kernels; (ii) NUMA‑aware memory allocation with enforced thread pinning; and (iii) aggressive loop unrolling combined with software prefetching. Consequently, three main gaps in the literature are filled: (i) open-source availability and reproducibility of all these low-level tricks in one place – ensuring that the improvement is not a one‑off. Many prior works were either proprietary or did not release their full code (e.g., HpQC, PAS), or they focused on a single aspect (e.g., PAS on single-node CPU, Intel-QS on distributed scaling) without covering the entire optimization spectrum. (ii) The lack of fine-grained architectural tuning. For example, most simulators do their optimizations internally and automatically; there is little ability to adjust, say, toggle on/off a specific optimization for comparison. Our work provides a more tunable environment, where one can study the effect of, e.g., enabling NUMA-aware allocation vs. not, or using AVX2 vs. AVX-512. This is highly valuable for scientific computing research, as it enables understanding the contribution of each optimization. It also aids in targeting different architectures – e.g., older CPUs without AVX-512 might benefit from different tuning than the latest servers – something an explicitly tunable simulator can accommodate. Finally, (iii) the focus on NUMA-aware execution is filling a clear gap. Prior open-source tools largely assumed a uniform shared-memory model and did not optimize for multiple NUMA regions, which can leave performance on the table in multi-socket machines. By implementing strategies such as allocating state subarrays per NUMA node and pinning threads, we can significantly reduce interconnect traffic.
@@
Paper Organization. Section II details our system architecture and optimization methodology—including our novel data structure layout (Section II-A), NUMA-aware optimization strategies (Sections II-B1 and II-B2), and various instruction-level tuning techniques (Section II-C). Section III presents our benchmark results and analysis, covering single- and two-qubit operations, QFT, and RQC performance. As demonstrated, integrating our newly designed data layout and these targeted optimizations yields performance gains comparable to those reported by HpQC and PAS, although not as pronounced as some of their most optimistic findings. Section IV discusses the scaling and portability of our framework. Finally, Section V concludes with a summary of our findings and a discussion of potential directions for future research.

### v2

@@ Abstract
Scalable classical simulation of quantum circuits is crucial for advancing quantum algorithm development and validating emerging hardware. This work focuses on performance enhancements through targeted low-level and NUMA-aware tuning on a single-node system, thereby not only advancing the efficiency of classical quantum simulations but also establishing a foundation for scalable, heterogeneous implementations that bridge toward noiseless quantum computing. Although few prior studies have reported similar hardware-level optimizations, such implementations have not been released as open-source software, limiting independent validation and further development. We introduce an open-source, high-performance extension to the QuEST state vector simulator that integrates state-of-the-art low-level and NUMA-aware optimizations for modern processors. Our approach emphasizes locality-aware computation and incorporates hardware-specific techniques including NUMA-aware memory allocation, thread pinning, AVX-512 vectorization, aggressive loop unrolling, and explicit memory prefetching. Experiments demonstrate substantial speedups–5.5-6.5 for single-qubit gate operations, 4.5 for two-qubit gates, 4 for Random Quantum Circuits (RQC), and 1.8 for the Quantum Fourier Transform (QFT). Algorithmic workloads further achieve 4.3–4.6 acceleration for Grover and 2.5 for Shor-like circuits. These results show that systematic, architecture-aware tuning can significantly extend the practical simulation capacity of classical quantum simulators on current hardware.
I Introduction
@@
Classical simulation of quantum circuits remains invaluable for validating algorithms, benchmarking hardware, and debugging control software while error-corrected quantum computers remain beyond current reach. In the full state‑vector simulation an -qubit state is represented by complex amplitudes, so both the memory footprint (16 bytes per amplitude in double-precision) and the number of floating‑point operations grow exponentially with . Even a modest 35‑qubit circuit already requires about 0.5 TiB of memory and billions of complex arithmetic operations, quickly exhausting today’s servers. Scaling beyond a few dozen qubits remains challenging because hardware and software bottlenecks dominate performance. Memory bandwidth and capacity are primary constraints: even on multi‑socket non‑uniform memory access (NUMA) servers, accessing a ‑element state vector saturates the memory hierarchy [1]. In distributed quantum-circuit simulation, the network communication required for qubit gates–whether single- or two-qubit–spanning different nodes introduces substantial overhead and often dominates the run time [1, 2]. At the software level, most simulators still under-use low‑level hardware features, leaving appreciable performance untapped and underscoring the need for architecture‑aware refinement.
@@
State-of-the-art and practice. To push the simulation limits, several high‑performance state‑vector simulators have emerged that exploit modern hardware architectures and parallelism [3]. QuEST [4], ProjectQ [5], Intel‑QS (qHiPSTER) [6], Qrack [7], Qulacs [8], Quantum++ [9], qsim [10], and more recently PennyLane Lightning [11] are representative state-of-the-art examples – summarized in Table I. Among many simulators exploiting multi‑threading (OpenMP), distributed Message Passing Interface (MPI) execution, or GPU off‑loading to enlarge the range of tractable problem sizes, some state vector simulators accelerate the core operation–applying gate matrices to the state vector–through explicit single-instruction-multiple-data (SIMD) parallelism. By operating on multiple amplitudes per CPU instruction, SIMD vectorization can significantly speed up state updates [4]. For example, the ProjectQ simulator explicitly supports Intel AVX vector instructions [5]. More recently, Google’s qsim is built with AVX2 and fused multiply-add (FMA) intrinsics, using fused gate kernels to exploit data-level parallelism [10]. Xanadu’s Lightning similarly employs explicit AVX-512 intrinsics, effectively unrolling loops to operate on 8 or 16 values at a time [11]. These approaches reduce instruction overhead by applying gate operations to multiple state vector entries concurrently. Academic simulators have reported substantial speedups from SIMD optimizations. The HpQC simulator used AVX2/FMA instructions to maximize SIMD utilization, combined with manual bit-level optimizations [12], yielding an average 2.20x speedup (GNU compiler) over QuEST on QFT benchmarks using a multi-core CPU. The authors attribute this gain largely to vectorized gate kernels and reduced instruction count. Another effort, PAS [13], introduced a hybrid vectorization scheme alongside other techniques, achieving multi-fold speedups versus QuEST.
@@
| Simulator | Vectorisation | Manual Unroll | Cache Blocking | OMP | Thread Pinning | NUMA | MPI | GPU | Open-Source |
| ProjectQ | AVX2 (comp.) | ✗ | ✓ | ✓ | — | ✗ | ✗ | ✗ | ✓ |
| qsim | AVX2 (intrin.) | ✗ | ✓ | ✓ | — | ✗ | ✗ | ✗ | ✓ |
| Lightning | AVX2/512 (intrin.) | ✓ | ✗ | ✓ | likely | ✗ | ✓ | ✓ | ✓ |
| HpQC/PAS | AVX2 (intrin.) | ✗ | ✗ | ✓ | — | ✗ | ✓ | ✗ | ✗ |
| Qulacs | AVX2 (intrin.) | ✗ | ✗ | ✓ | — | ✗ | ✗ | ✓ | ✓ |
| Qiskit Aer | AVX2 (comp.) | ✗ | ✗ | ✓ | — | ✗ | ✗ | ✓ | ✓ |
| Intel-QS | AVX512 (comp.) | ✗ | ✗ | ✓ | likely | ✗ | ✓ | ✗ | ✓ |
| QuEST | AVX2 (comp.) | ✗ | ✗ | ✓ | — | ✗ | ✓ | ✓ | ✓ |
| NWQ-Sim | AVX2 (comp.) | ✗ | ✗ | ✓ | — | ✗ | ✓ | ✓ | ✓ |
“Aligned Memory”, not included, is largely implicit since most simulators use 64-byte alignment for AVX-512 operations. “Thread Pinning” and “NUMA-Aware Allocation” are seldom addressed explicitly; in most cases, thread affinity follows the default OpenMP scheduler unless the user enforces binding via environment variables such as OMP_PROC_BIND or OMP_PLACES. PennyLane Lightning and Intel-QS are inferred to support such affinity through OpenMP configuration rather than hard-coded pinning. comp. = compiler auto-vectorisation; intrin. = handwritten intrinsics. likely = affinity supported or recommended (e.g., via OpenMP settings) but not explicitly enforced in source code. “—” = no explicit thread-affinity or pinning support (default OpenMP scheduling). ✓/✗ = feature explicitly present or absent according to publication, documentation, or verified source inspection.
@@
To fully leverage SIMD, simulators ensure proper memory alignment of the state vector (often 32- or 64-byte aligned for AVX2/512). Although often not explicitly discussed in publications, memory alignment is a standard practice to avoid penalties when loading vectors into registers. Many simulators also focus on cache-friendly memory access patterns. A common technique is cache blocking–structuring computations to maximize reuse of data in fast caches before accessing slower memory. In quantum-circuit terms, this translates to fusing consecutive gate operations so that the state vector is traversed fewer times. At a higher level, circuits can also be rearranged to minimize communication, effectively cache-blocking at the node level [14, 15]. ProjectQ and Google’s qsim use such gate fusion to improve locality. By multiplying small gate matrices into a larger effective matrix, the simulator can apply a batch of operations with one pass over the state, thus improving cache utilization and reducing memory bandwidth pressure. Beyond fusion, recent research introduced novel data layouts and filters to minimize unnecessary memory traffic. PAS describes a “memory access filtering” method, which avoids touching state vector regions that are unaffected by a given gate. Likewise, PAS’s use of fast bitwise operations optimizes the index calculations for mapping qubit operations to memory locations, replacing expensive arithmetic or branch-heavy code with efficient bit-manipulation. The HpQC tool also emphasized a custom data structure to exploit spatial locality in the cache. In practice, these low-level memory optimizations yielded measurable speedups: e.g., HpQC reported 1.74 (GNU compiler) speed improvement on RQCs vs. baseline QuEST by cutting down redundant memory accesses.
@@
Today’s Hardware. Modern CPUs feature dozens of cores, and all state vector simulators employ multithreading to split the amplitude space across cores. OpenMP is commonly employed to parallelize the innermost loops of state updates. QuEST and other simulators leverage multi-core parallelism by dividing the state vector into chunks processed concurrently by threads. This yields near-linear speedup with core count [4] until memory bandwidth becomes the bottleneck [16]. However, multithreading on large shared-memory systems introduces challenges related to thread affinity and load balancing. If threads are not carefully pinned or if the work is unevenly distributed, performance can suffer. Thread pinning (assigning each thread to a specific core) is often required to prevent operating system (OS) scheduler overhead or migration between cores. Additionally, on multi-socket NUMA nodes, pinning helps ensure threads access memory local to their socket. While most quantum simulator papers do not explicitly describe their thread affinity policies, the need for such control is well recognized in the community [17].
@@
On NUMA architectures, memory is physically partitioned across sockets, and accessing remote memory through the interconnect (e.g., QPI/UPI) occurs transparently, i.e., without explicit communication, but incurs higher latency and lower bandwidth. It is therefore crucial for large state vector simulations to be NUMA-aware. A straightforward strategy is to partition the state vector across NUMA nodes and pin threads such that each domain primarily works on its local portion of the data. In the ideal case, each thread only accesses the segment of the state vector residing in its local node’s memory [17]. This avoids the costly scenario where a thread continuously fetches data from another socket’s memory. When thread affinity and memory placement are aligned, memory bandwidth scales with the number of NUMA nodes instead of being bottlenecked by slower cross-socket memory access [18]. Despite its importance, NUMA-aware optimization has not always been a primary focus in quantum simulator research. Many open-source simulators rely on the OS’s default first-touch policy for memory allocation (as implemented in Linux [19]) but do not report additional NUMA-specific strategies in their publications. We found very little to no explicit analysis of NUMA effects in simulator papers, indicating a clear gap that our work aims to address. A recent IBM patent [17] describes how a naïve parallelization of a CNOT gate can lead to one socket’s thread repeatedly accessing another socket’s memory, an imbalanced situation, and proposes reordering qubit indices or replicating data to balance memory accesses. The fact that such patent-level solutions exist suggests that while NUMA issues are recognized, they have not been thoroughly explored in open literature.
Contributions. Motivated by these observations, and building on the robust foundation of QuEST v3.7.0111Our baseline throughout this work is the default QuEST v3.7.0 build compiled with AVX2-enabled compiler optimizations (e.g., -O3 -mavx2) but without explicit intrinsics, FMA operations, or NUMA/thread placement controls. In preparing this draft, we noted that QuEST v4.0 was released, introducing several software architectural enhancements–including revised data structures that partially overlap with our improvements–but it does not incorporate explicit hand-written AVX-512 intrinsics, cache-aware loop unrolling, prefetching, and NUMA pinning/partitioning as presented in this work. Details are discussed in Section V. we present the first open-source, comprehensive evaluation of low-level and NUMA-aware optimizations for classical quantum simulation, providing a suite of single‑node enhancements forming a tunable foundation for future distributed and heterogeneous (multi-node and GPU) scaling. Our work consolidates and generalizes multiple architecture-specific techniques–previously studied only in isolation or without public implementations, see Table I–into a single reproducible framework. Specifically, this paper contributes:
(1) Open-source and reproducible low-level optimization framework: a publicly available extension to QuEST integrating, for the first time, all key architectural optimizations–including explicit AVX-512 intrinsics, cache-aware loop unrolling, software prefetching, and NUMA-aware thread pinning–within a unified, benchmarked codebase. This enables transparent evaluation and reuse by the research community, ensuring that the improvement is not a one‑off. Many prior works were either proprietary or did not release their full code (e.g., HpQC, PAS), or they focused on a single aspect (e.g., PAS on single-node CPU, Intel-QS on distributed scaling) without covering the entire optimization spectrum.
(2) Fine-grained architectural tunability: unlike most simulators that apply compiler-driven or opaque internal optimizations, our implementation exposes each optimization (e.g., AVX2 vs. AVX-512, loop-unrolling factor, prefetching distance, NUMA-aware vs. default memory allocation) as a configurable module. This tunable design enables researchers to systematically examine the impact of individual optimizations and understand their individual contribution to overall performance. Such flexibility is particularly valuable for scientific computing, as it supports reproducible and transparent performance evaluation across diverse hardware architectures. It also allows the simulator to be easily adapted to different processor generations; for example, older CPUs without AVX-512 may benefit from smaller unrolling factors or alternative memory-placement strategies, which our framework can selectively enable or disable.
(3) Explicit NUMA-aware execution: we provide a detailed NUMA-partitioned memory model combined with enforced thread–memory affinity and locality-sensitive scheduling. This minimizes inter-socket memory traffic and scales memory bandwidth nearly linearly with the number of NUMA nodes–addressing a critical gap left by previous open-source simulators that assumed uniform memory access, which can leave performance on the table in multi-socket machines.
(4) Comprehensive performance study: evaluating low-level optimizations across six representative workloads; single-qubit, two-qubit, Random Quantum Circuits (RQC), Quantum Fourier Transform (QFT), Grover, and Shor-like circuits, providing a holistic view of how architectural tuning affects distinct quantum workloads.
Our focus in this work is on mathematical, kernel-level architectural optimizations within QuEST rather than on full cross-simulator benchmarking. We systematically examine how different low-level optimizations affect performance within a single software framework. Prior studies, including the original QuEST paper [4] and recent Queen quantum simulator [20], have consistently shown QuEST to be competitive or superior to simulators such as Qiskit Aer, Intel-QS, cuQuantum, and qsim across a wide range of architectures and workloads. This supports our choice of QuEST as a representative high-performance baseline. We specifically target node-local memory bottlenecks–an often overlooked yet critical aspect of quantum simulation performance. Similar to PAS, HpQC, and Qulacs, we begin with single-node optimization as a necessary foundation for scalable multi-node execution. We prioritize deep, low-level kernel tuning over early MPI distribution, since premature scaling without resolving intra-node memory and threading bottlenecks can degrade overall performance. By first addressing locality and vectorization challenges, we establish a robust baseline for subsequent distributed extensions, which are under active development and will be presented in a follow-up study.
@@
Paper Organization. Section II details our system architecture and optimization methodology—including our novel data structure layout (Section II-A), NUMA-aware optimization strategies (Sections II-B1 and II-B2), and various instruction-level tuning techniques (Section II-C). Section III presents our benchmark results and analysis, covering the performance of single- and two-qubit gate operations, QFT, RQC, Grover, and a Shor-like periodicity estimation circuit. As demonstrated, integrating our newly designed data layout and these targeted optimizations yields performance gains comparable to those reported by HpQC and PAS, although not as pronounced as some of their most optimistic findings. Section IV discusses the scaling and portability of our framework. Finally, Section V concludes with a summary of our findings and a discussion of potential directions for future research.